### v3

@@ 6 Conclusions
where the last equality follows by the outer semicontinuity of the limiting subdifferential [57, Prop. 8.7] with the continuity of $J$ at $\widebar{\bm{z}}$.
To show the first inclusion in (101), let $\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$, i.e., there exists a sequence $(\bm{v}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\bm{v}_{n}\in A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$ such that its subsequence $(\bm{v}_{m(l)})_{l=1}^{\infty}$ converges to $\bm{v}$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, there exists a sequence $(\bm{u}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ with $\bm{u}_{n}\in\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ satisfying $\bm{v}_{n}=A_{n}(\bm{u}_{n})$. From Fact A.3, the local Lipschitz continuity of $J$ at $\widebar{\bm{z}}$ ensures the boundedness of $\bigcup_{n=n_{0}}^{\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ with a sufficiently large $n_{0}\in\mathbb{N}$. This implies the boundedness of $(\bm{u}_{n})_{n=1}^{\infty}$, thereby its subsequence $(\bm{u}_{m(l)})_{l=1}^{\infty}$ converges to some $\bm{u}\in\mathcal{Z}$ (by passing to further subsequence if necessary), i.e., $\bm{u}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$. We get $\bm{v}=\lim_{l\to\infty}A\bm{u}_{m(l)}$ by
@@
For $n\in\mathbb{N}$, let $\widebar{\bm{z}}_{n}\coloneqq\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))$ and $A_{n}\coloneqq(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}$. Then, we have $\lim_{n\to\infty}A_{n}=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}\eqqcolon A$ from the continuous differentiability of $\mathcal{F}$. Recall that $J$ is locally Lipschitz continuous [^19] Since every finite convex function is locally Lipschitz continuous [57, Exm. 9.14], $J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ and $\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ are locally Lipschitz continuous due to the weak convexity of $J$. Then, $J=(J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})+(-\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})$ is locally Lipschitz continuous because the sum of locally Lipschitz continuous functions is locally Lipschitz continuous [57, Exe. 9.8]. , and thus $\partial_{\rm L}J(\bm{z})\neq\emptyset$ for all $\bm{z}\in\mathcal{Z}$, by Fact A.3, i.e., $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$. From $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$, $\lim_{n\to\infty}\mathcal{F}(\bm{y}_{n})=\mathcal{F}(\widebar{\bm{y}})$, and $\mu_{n}\searrow 0$, Fact 2.7 (a) implies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))=P_{\overline{\mathrm{dom}(\partial_{\rm L}J)}}(\mathcal{F}(\widebar{\bm{y}}))=\mathcal{F}(\widebar{\bm{y}})\eqqcolon\widebar{\bm{z}}$. From the local Lipschitz continuity of $J$, Lemma C.1 with $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$ and $\lim_{n\to\infty}A_{n}=A$ yields
@@
where the first equality follows by the fact $\liminf_{n\to\infty}d(\bm{v},E_{n})=d\left(\bm{v},\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\right)$ for $\bm{v}\in\mathcal{Y}$ and a given sequence $(E_{n})_{n=1}^{\infty}$ of subsets $E_{n}\subset\mathcal{Y}$ (see [57, Exe. 4.8]). ∎
Appendix D Proof of Corollary 5.1
@@
Fact D.1 ([54, Lemma A.4 and proof in Proposition 2.9]).
@@
Lemma D.2 (Boundedness and Lipschitz continuity of $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$).
@@
(Proof of (119)) Let $\bm{V},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact D.1 (d), we have
@@
(Proof of (120)) Let $\bm{V}_{1},\bm{V}_{2},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact D.1 (d), we have
@@
(a) By noting that $\bm{\varphi}_{\bm{S}}^{\rm Cay}:Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a diffeomorphism [54, Prop. 2.2], Corollary 3.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ completes the proof [^21] Although $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is not surjective onto ${\rm St}(p,N)$, the same statement in Corollary 3.4 follows (see Remark 3.5). (see also the footnote7 in Corollary 3.4).
(b) To invoke Proposition 4.2, we check below the conditions (A1)-(A5) in Proposition 4.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. Since ${\rm St}(p,N)$ is compact and $\mathfrak{S}$ is twice continuously differentiable, (A1)-(A3) are automatically satisfied (see Remark 4.3 (a)). For (A4)-(A5), see Lemma D.2. Therefore, by Proposition 4.2 (b), Assumption 4.7 (a) is satisfied with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. ∎

### v4

@@ A.2 Penalty-based strategy
where the last equality follows by the outer semicontinuity of the limiting subdifferential [69, Prop. 8.7] with the continuity of $J$ at $\widebar{\bm{z}}$.
To show the first inclusion in (106), let $\bm{v}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$, i.e., there exists a sequence $(\bm{v}_{n})_{n=1}^{\infty}\subset\mathcal{Y}$ with $\bm{v}_{n}\in A_{n}(\partial_{\rm L}J(\widebar{\bm{z}}_{n}))$ such that its subsequence $(\bm{v}_{m(l)})_{l=1}^{\infty}$ converges to $\bm{v}$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, there exists a sequence $(\bm{u}_{n})_{n=1}^{\infty}\subset\mathcal{Z}$ with $\bm{u}_{n}\in\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ satisfying $\bm{v}_{n}=A_{n}(\bm{u}_{n})$. From Fact B.3, the local Lipschitz continuity of $J$ at $\widebar{\bm{z}}$ ensures the boundedness of $\bigcup_{n=n_{0}}^{\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$ with a sufficiently large $n_{0}\in\mathbb{N}$. This implies the boundedness of $(\bm{u}_{n})_{n=1}^{\infty}$, thereby its subsequence $(\bm{u}_{m(l)})_{l=1}^{\infty}$ converges to some $\bm{u}\in\mathcal{Z}$ (by passing to further subsequence if necessary), i.e., $\bm{u}\in\mathop{\mathrm{Lim~sup}}_{n\to\infty}\partial_{\rm L}J(\widebar{\bm{z}}_{n})$. We get $\bm{v}=\lim_{l\to\infty}A\bm{u}_{m(l)}$ by
@@
For $n\in\mathbb{N}$, let $\widebar{\bm{z}}_{n}\coloneqq\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))$ and $A_{n}\coloneqq(\mathrm{D}\mathcal{F}(\bm{y}_{n}))^{*}$. Then, we have $\lim_{n\to\infty}A_{n}=(\mathrm{D}\mathcal{F}(\widebar{\bm{y}}))^{*}\eqqcolon A$ from the continuous differentiability of $\mathcal{F}$. Recall that $J$ is locally Lipschitz continuous [^22] Since every finite convex function is locally Lipschitz continuous [69, Exm. 9.14], $J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ and $\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ are locally Lipschitz continuous due to the weak convexity of $J$. Then, $J=(J+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})+(-\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2})$ is locally Lipschitz continuous because the sum of locally Lipschitz continuous functions is locally Lipschitz continuous [69, Exe. 9.8]. , and thus $\partial_{\rm L}J(\bm{z})\neq\emptyset$ for all $\bm{z}\in\mathcal{Z}$, by Fact B.3, i.e., $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$. From $\mathrm{dom}(\partial_{\rm L}J)=\mathcal{Z}$, $\lim_{n\to\infty}\mathcal{F}(\bm{y}_{n})=\mathcal{F}(\widebar{\bm{y}})$, and $\mu_{n}\searrow 0$, Fact 2.7 (a) implies $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\lim_{n\to\infty}\mathrm{prox}_{\mu_{n}J}(\mathcal{F}(\bm{y}_{n}))=P_{\overline{\mathrm{dom}(\partial_{\rm L}J)}}(\mathcal{F}(\widebar{\bm{y}}))=\mathcal{F}(\widebar{\bm{y}})\eqqcolon\widebar{\bm{z}}$. From the local Lipschitz continuity of $J$, Lemma D.1 with $\lim_{n\to\infty}\widebar{\bm{z}}_{n}=\widebar{\bm{z}}$ and $\lim_{n\to\infty}A_{n}=A$ yields
@@
where the first equality follows by the fact $\liminf_{n\to\infty}d(\bm{v},E_{n})=d\left(\bm{v},\mathop{\mathrm{Lim~sup}}_{n\to\infty}E_{n}\right)$ for $\bm{v}\in\mathcal{Y}$ and a given sequence $(E_{n})_{n=1}^{\infty}$ of subsets $E_{n}\subset\mathcal{Y}$ (see [69, Exe. 4.8]). ∎
Appendix E Proof of Corollary 5.1
@@
Fact E.1 ([48, Lemma A.4 and proof in Proposition 2.9]).
@@
Lemma E.2 (Boundedness and Lipschitz continuity of $\mathrm{D}\bm{\varphi}_{\bm{S}}^{\rm Cay}(\bm{V})$).
@@
(Proof of (124)) Let $\bm{V},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact E.1 (d), we have
@@
(Proof of (125)) Let $\bm{V}_{1},\bm{V}_{2},\bm{D}\in Q_{N,p}$ such that $\left\lVert\bm{D}\right\rVert_{F}\leq 1$. By Fact E.1 (d), we have
@@
(a) By noting that $\bm{\varphi}_{\bm{S}}^{\rm Cay}:Q_{N,p}\to{\rm St}(p,N)\setminus E_{N,p}(\bm{S})$ is a diffeomorphism [48, Prop. 2.2], Corollary 3.4 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$ completes the proof [^24] Although $\bm{\varphi}_{\bm{S}}^{\rm Cay}$ is not surjective onto ${\rm St}(p,N)$, the same statement in Corollary 3.4 follows (see Remark 3.5). (see also the footnote7 in Corollary 3.4).
(b) To invoke Proposition 4.2, we check below the conditions (A1)-(A5) in Proposition 4.2 with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. Since ${\rm St}(p,N)$ is compact and $\mathfrak{S}$ is twice continuously differentiable, (A1)-(A3) are automatically satisfied (see Remark 4.3 (a)). For (A4)-(A5), see Lemma E.2. Therefore, by Proposition 4.2 (b), Assumption 4.7 (a) is satisfied with $C\coloneqq{\rm St}(p,N)$ and $\bm{\varphi}\coloneqq\bm{\varphi}_{\bm{S}}^{\rm Cay}$. ∎